OpenStudy (anonymous):

Please don't answer this unless you know for a fact, cuz lately people have been tellin me the wrong answer, anyways A community garden group is giving boxes of plants to local gardeners. The people who assembled the boxes randomly picked items from two types of vegetable plants, 4 kinds of trees, 5 kinds of flowers, and 5 types of herbs. Find the probability that Peter will pick a box with a pea plant, a white ash tree, a daylily, and a tarragon herb plant.

OpenStudy (studygurl14):

2 types of vegetables 4 kinds of trees 5 kinds of flower 5 types of herbs You want the probability one of each plant...so... \[\frac{1}{2}\times \frac{1}{4} \times \frac{1}{5} \times \frac {1}{5} = ?\]

OpenStudy (anonymous):

so 1/200?

OpenStudy (anonymous):

ok got it thx

OpenStudy (studygurl14):

yep yw